1. ABC, Inc. is expected to pay an annual dividend per share of $2.80, and investors require a rate of return on S&P500 index is 12%, with the T-bill rate at 6%. What should be the current share price of ABC's stock if ABC's beta is 1.7, with a growth rate of 5%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

2.

If firm A has a higher plowback ratio than firm B, then firm A will tend to have ___________ P/E ratio than firm B, other things being equal.

Multiple Choice

  • a lower

  • a higher

  • an unchanged

  • The answer cannot be determined from the information given.

3.

A preferred stock will be worth ______________ today if it is expected to pay a fixed amount of $6 dividend per share every year, and investors require a return of 7% on this preferred stock.

Multiple Choice

  • $65.50

  • $85.71

  • $114.29

  • $33.50
Answer #1

Hi

As per policy we will solve only 1 question here.

Here annual dividend D1 = $2.80

S&P return (market return) rm = 12%

T bill rate (risk free rate) rf = 6%

beta = 1.7

As per Capital Asset pricing model

required rate of return of stock r= rf + beta*(rm-rf)

= 6 +1.7*(12-6)

=6+ 1.7*6

= 16.2%

growth rate g = 5%

As per dividend discount model

Current Stock Price = D1/(r-g)

= 2.8/(16.2%- 5%)

= $25

Thanks